A random sample of five policyholders is taken from a group of 10 comprising 4men and 6women. Let x represent the number of females in the sample. Calculate the expected value and variance of the number of female policyholders


1
Expert's answer
2021-12-02T06:10:25-0500
"\dbinom{10}{5}=\dfrac{10!}{5!(10-5)!}=\dfrac{10(9)(8)(7)(6)}{120}=252"


"P(X=0)=0"

"P(X=1)=\dfrac{\dbinom{6}{1}\dbinom{4}{4}}{\dbinom{10}{5}}=\dfrac{6}{252}"

"P(X=2)=\dfrac{\dbinom{6}{2}\dbinom{4}{3}}{\dbinom{10}{5}}=\dfrac{15(4)}{252}=\dfrac{60}{252}"

"P(X=3)=\dfrac{\dbinom{6}{3}\dbinom{4}{2}}{\dbinom{10}{5}}=\dfrac{20(6)}{252}=\dfrac{120}{252}"

"P(X=4)=\dfrac{\dbinom{6}{4}\dbinom{4}{1}}{\dbinom{10}{5}}=\dfrac{15(4)}{252}=\dfrac{60}{252}"

"P(X=5)=\dfrac{\dbinom{6}{5}\dbinom{4}{0}}{\dbinom{10}{5}}=\dfrac{6}{252}"

"\def\arraystretch{1.5}
 \begin{array}{c:c}
 x & 0 & 1 & 2 & 3 & 4 & 5 \\ \hline
 p(x) & 0 & 1/42 & 10/42 & 20/42 & 10/42 & 1/42 \\
 
\end{array}"


"E(X)=0(0)+\dfrac{1}{42}(1)+\dfrac{10}{42}(2)+\dfrac{20}{42}(3)"

"+\dfrac{10}{42}(4)+\dfrac{1}{42}(5)=3"


"E(X^2)=0(0)+\dfrac{1}{42}(1)^2+\dfrac{10}{42}(2)^2+\dfrac{20}{42}(3)^2"

"+\dfrac{10}{42}(4)^2+\dfrac{1}{42}(5)^2=\dfrac{29}{3}"

"Var(X)=E(X^2)-[E(X)]^2"

"=\dfrac{29}{3}-(3)^2=\dfrac{2}{3}"